OpenStudy (scarlettfarra2000):

Will someone check and help me with the next part Multiply and simplify by factoring

OpenStudy (scarlettfarra2000):

\[\sqrt{14}\times \sqrt{6}\]

OpenStudy (scarlettfarra2000):

The first step is to combine the radicals using the product rule radicals \[\sqrt{a}\times \sqrt{b}=\sqrt{ab}\] \[\sqrt{14}\times \sqrt{6}=\sqrt{14\times6}=\sqrt{84}\]

OpenStudy (photon336):

yep

OpenStudy (scarlettfarra2000):

Next,factor the radicand so that one factor is a perfect square \[\sqrt{84}=?\] I know what the factors of 84 is but which two do I use? 1,2,3,4,6,7,14,21,28,42,84

OpenStudy (scarlettfarra2000):

I was thinking \[\sqrt{21}\times \sqrt{4}\]

OpenStudy (photon336):

excellent

OpenStudy (photon336):

that's right, so what would you end up getting?

OpenStudy (scarlettfarra2000):

now factor into two radicals \[\sqrt{21\times4}=\sqrt{21}\times \sqrt{4}\] finally, simplify each radical, if possible \[\sqrt{21}\times \sqrt{4}=7\sqrt{4}\]

OpenStudy (photon336):

remember \[2\sqrt{21}\]

OpenStudy (photon336):

the reason for this is that there aren't any perfect squares that go into 21

OpenStudy (scarlettfarra2000):

Thank you @Photon336